OpenStudy (anonymous):

Solve the following system of equations for x: 5x + 6 = 4x + y + 16 3(x - 2) – y = 2x - 3y + 5 Just to check. When I did this, I got x = 13

OpenStudy (anonymous):

The following is not your normal class room solution. Eliminate is a Mathematica function that in the case of the expression below will eliminate y. Eliminate[ {6+5 x=16+4 x+y,3 (-2+x)-y=5+2 x-3 y}, y ] 3x = 31\[x=\frac{31}{3} \]

OpenStudy (anonymous):

\[\left\{x=\frac{31}{3},y=\frac{1}{3}\right\} \] is the solution to the equations.

OpenStudy (anonymous):

I don't really understand how you got those answers.

OpenStudy (anonymous):

Class room method: Solve the first equation for y and simplify. y = x - 10 Now replace each y in the second equation with (x-10) and simplify.

OpenStudy (anonymous):

Ohh. Okay, that makes sense.

OpenStudy (anonymous):

{10 + 3 (-2 + x) - x = 5 - 3 (-10 + x) + 2 x}\[4+2 x=35-x\]3x = 31 x = 31/3

OpenStudy (anonymous):

Thank you for the medal.